OpenStudy (anonymous):

show steps to simpliying this ratio: x^2+8x / x^2+10x+16

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x^2+8x}{x^2+10x+16}\] \[\Large \frac{x(x+8)}{x^2+10x+16}\] \[\Large \frac{x(x+8)}{(x+8)(x+2)}\] \[\Large \frac{x\cancel{(x+8)}}{\cancel{(x+8)}(x+2)}\] \[\Large \frac{x}{x+2}\] So \[\Large \frac{x^2+8x}{x^2+10x+16}=\frac{x}{x+2}\]